Jupiter Company's budget for the coming year estimates monthly sales in units for the first five months of the year as follows: April, 35,000; May, 38,000; June 40,000; July, 39,000; and August, 38,000.

The ending inventory of finished units available for sale is to be maintained at 25% of the next month's estimated sales. On April 1, there were 12,000 units on hand. There is no work in process at the end of any month. Each unit of finished product requires the use of five units of materials.

Materials are to be carried in inventory in an amount equal to one-third of the expected production of the coming month. Beginning inventory on April 1 of the current year was 54,167 units of material.

a. Prepare a production budget (in units) for the first four months of the period. Round amounts

to the nearest unit.

b. Prepare a materials purchases budget (in units of materials) for the second quarter of the

coming year. Round amounts to the nearest unit.
Answer #1

Answer-a)-

JUPITER COMPANY
PRODUCTION BUDGET
FOR THE FIRST FOUR MONTHS
Particulars April May June July Total
Estimated sales (a) 35000 38000 40000 39000 152000
Add:- Desired ending finished goods inventory (b) 38000*25%=9500 40000*25%=10000 39000*25%=9750 38000*25%=9500 38750
25% of the next month's sales needs
Less:-Opening finished goods inventory (c) 12000 9500 10000 9750 41250
Number of units to be produced (d=a+b-c) 32500 38500 39750 38750 149500

b)-

JUPITER COMPANY
DIRECT MATERIALS PURCHASES BUDGET
FOR THE PERIOD ENDED SECOND QUARTER
Particulars APRIL MAY JUNE Year
Required production in units of finished goods (a) 32500 38500 39750 110750
Units of raw material required per units of finished goods (b) 5 5 5 5
Units of raw material needed to meet production a*b=(c) 162500 192500 198750 553750
Add:- Desired units of ending raw material inventory 192500*1/3=64167 198750*1/3=66250 193750*1/3=64583 195000
Total units of raw material needed 226667 258750 263333 748750
Less:-Units of opening raw material inventory 54167 67167 66250 187584
Units of raw material to be purchased (d) 172500 191583 197083 561166